OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\huge \frac { \sqrt { 6 } }{ { \sqrt { 2 } +\sqrt { 3 } } } +\frac { 3 }{ { \sqrt { 6 } +\sqrt { 3 } } } -\frac { 4\sqrt { 3 } }{ { \sqrt { 6 } +\sqrt { 2 } } } \]

OpenStudy (anonymous):

@vbfn

OpenStudy (anonymous):

@vbfn

OpenStudy (anonymous):

hmmm.......

OpenStudy (anonymous):

make same denominator

OpenStudy (anonymous):

@satellite73 help

OpenStudy (anonymous):

isn't that just rationalizing denum??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

yo helpis

OpenStudy (anonymous):

\[4\sqrt{2}-3\sqrt{3}\]

OpenStudy (anonymous):

first term is \(3\sqrt{2}-2\sqrt{3}\)

OpenStudy (anonymous):

hmm.. cant get you

OpenStudy (anonymous):

second term is \(\sqrt{6}-\sqrt{3}\)

OpenStudy (anonymous):

you have to multiply by the conjugate 3 times, i don't see a snap way to do it other than than then combine like terms

OpenStudy (anonymous):

hmm......

OpenStudy (anonymous):

got that next ....???

OpenStudy (anonymous):

\[\frac{\sqrt{6}}{\sqrt{2}+\sqrt{3}}\times \frac{\sqrt{2}-\sqrt{3}}{\sqrt{2}-\sqrt{3}}\] \[=\frac{\sqrt{6}(\sqrt{2}-\sqrt{3})}{2-3}\] \[=\frac{2\sqrt{3}-3\sqrt{2}}{-1}\] \[=3\sqrt{2}-2\sqrt{3}\] etc etc etc

OpenStudy (anonymous):

that is the first term. second one and third do the same way

OpenStudy (anonymous):

okay thanks satellite

OpenStudy (anonymous):

second

OpenStudy (anonymous):